OpenStudy (anonymous):

Algebra 2 help: Solve. Check for extraneous solutions. (√5-x) - √x = 1 Also, "5-x" is all under a radicand I have an idea of what to do to solve it, but if someone could walk me through it that would be awesome (final answer is 1 [according to key])

OpenStudy (saifoo.khan):

sqrt(5-x) = 1 + sqrt(x) take square on both sides: 5-x = (1+sqrt(x)) ^2 5 - x = (1+sqrtx)(1+sqrtx) 5 - x = 1 + 2sqrt(x) + x 5 - 2x - 1 = 2sqrt x (4 - 2x) / 2 = sqrt x 2 - x = sqrt x Taking square on both sides: (2-x)^2 = (sqrt x)^2 (2-x)^2 = x

OpenStudy (saifoo.khan):

Now solve this and you'll get 2 valuse..

OpenStudy (saifoo.khan):

values*